OpenStudy (iwanttogotostanford):

HELP ASAP PLEASE Question:

OpenStudy (iwanttogotostanford):

@explorerhills @sweetburger

OpenStudy (iwanttogotostanford):

@Photon336

OpenStudy (iwanttogotostanford):

@sweetburger

OpenStudy (iwanttogotostanford):

please help!!

OpenStudy (photon336):

Wait so all three of those have to be roots?

OpenStudy (photon336):

@nincompoop

OpenStudy (agent0smith):

Remember the conjugate zeroes theorem? If a+bi is a zero, then a-bi is also a zero. So they gave you 8, -14, and 3 + 9i... but there's one more, what is it?

OpenStudy (agent0smith):

Once you realize that fourth zero, just plug all the zeroes in like so: \[\large (x- [zero])(x- [zero])(x- [zero])(x- [zero])\]replacing the [zero] with the given zeroes, and simplify.

OpenStudy (iwanttogotostanford):

so how would I get my answer though

OpenStudy (iwanttogotostanford):

can you just tell me the steps please, I am writing a study guide and need a solid example:-)

OpenStudy (iwanttogotostanford):

(this one^) please

OpenStudy (iwanttogotostanford):

@agent0smith

OpenStudy (agent0smith):

Go back and answer my question, first.

OpenStudy (iwanttogotostanford):

@agent0smith the other 0 is 2

OpenStudy (agent0smith):

No it isn't. Remember the conjugate zeroes theorem? If a+bi is a zero, then a-bi is also a zero. So they gave you 8, -14, and 3 + 9i... but there's one more, what is it? Eg. if a zero was 3-5i, another zero must be 3+5i.

OpenStudy (iwanttogotostanford):

so the other is 3-9i

OpenStudy (agent0smith):

Yes.

OpenStudy (iwanttogotostanford):

so what would i do next

OpenStudy (agent0smith):

I gave the next step already.

OpenStudy (iwanttogotostanford):

and then thats it once i do that?

OpenStudy (iwanttogotostanford):

or not distribute, factor out

OpenStudy (iwanttogotostanford):

@agent0smith

OpenStudy (agent0smith):

That's the long hard part, so far you've only done the easy part (x-8)(x+14)(x-3+9i)(x-3-9i) Start by expanding the (x-8)(x+14) Then expand the (x-3+9i)(x-3-9i)

OpenStudy (iwanttogotostanford):

x^4 - 11x^3 + 72x^2 - 606x + 10,080

OpenStudy (agent0smith):

Idk maybe. I assume you just plugged that into wolfram or something.

OpenStudy (iwanttogotostanford):

ok

OpenStudy (iwanttogotostanford):

thanks you're a big help. Lol i did it and thats what I got!! sheesh

OpenStudy (iwanttogotostanford):

never mind its not right, i redid it and its the last one. Thanks fort the help

OpenStudy (agent0smith):

idk what the options are but scroll down to see the correct expression here https://www.wolframalpha.com/input/?i=(x-8)(x%2B14)(x-3%2B9i)(x-3-9i)

OpenStudy (mww):

An easy way to expand the pesky ones (z' is conjugate of z): \[(x - z)( x - z') = x^2 - xz'-xz +zz' =x^2 -x(z+z')+zz' \] \[= x^2 - 2 \mathbb{R}e(z) x +|z|^2\]